OpenStudy (anonymous):

verify: sin(x+y)-sin(x-y)=2cosxsiny

OpenStudy (anonymous):

cos(x+y)=cosxacosy-sinxsiny cos(x-y)=cosxcosy+sinxsiny addition produces 2cosxcosy

OpenStudy (lgbasallote):

take the left hand side...do you know the formula for sin(x+y) and sin(x-y)?

OpenStudy (anonymous):

Its an identity.

OpenStudy (anonymous):

i think so but i'm not sure

OpenStudy (lgbasallote):

sin(x+y) = sinxcosy + sinycosx sin(x-y) = sinxcosy - sinycosx can you do it from here?

OpenStudy (anonymous):

it looks familiar but i can't figure it out. is it one of the cosine identities?

OpenStudy (lgbasallote):

no...you substitute those into your problem sin(x+y) - sin (x-y) = sinxcosy + cosxsiny - (sinxcosy - cosxsiny)

OpenStudy (anonymous):

i'm still confused

OpenStudy (lgbasallote):

with what?

OpenStudy (anonymous):

how to do all of this in general

OpenStudy (lgbasallote):

basically you memorize the identities...

OpenStudy (lgbasallote):

i did not do anything fancy...all i did was substitute identities

OpenStudy (anonymous):

oh ok

OpenStudy (lgbasallote):

so regarding this problem...you're good?

OpenStudy (anonymous):

not really...

OpenStudy (lgbasallote):

sinxcosy + cosxsiny - (sinxcosy - cosxsiny) sinxcosy = cosxsiny - sinxcosy + cosxsiny

OpenStudy (lgbasallote):

that's a + not an =

OpenStudy (anonymous):

i don't understand what you did

OpenStudy (lgbasallote):

i distributed the minus sign

OpenStudy (anonymous):

oh ok i see

OpenStudy (lgbasallote):

so you're good now?

OpenStudy (anonymous):

yes thanks ^^